OpenStudy (anonymous):

Which number is divisible by 4? will medal and fan no direct answers tho thank you c: A. 38 B. 74 C. 116 D. 134

OpenStudy (anonymous):

Okay I can help you

OpenStudy (anonymous):

38/4 = 9.5 74/4 = 18.5 116/4 = 29 134/4 = 33.5

OpenStudy (anonymous):

so c?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

no problem...any other questions?